# AI-Driven Guided Response for Security Operation Centers with Microsoft Copilot for Security

### v3

@@ Abstract.
Security operation centers contend with a constant stream of security incidents, ranging from straightforward to highly complex. To address this, we developed Microsoft Copilot for Security Guided Response (CGR), an industry-scale ML architecture that guides security analysts across three key tasks—(1) investigation, providing essential historical context by identifying similar incidents; (2) triaging to ascertain the nature of the incident—whether it is a true positive, false positive, or benign positive; and (3) remediation, recommending tailored containment actions. CGR is integrated into the Microsoft Defender XDR product and deployed worldwide, generating millions of recommendations across thousands of customers. Our extensive evaluation, incorporating internal evaluation, collaboration with security experts, and customer feedback, demonstrates that CGR delivers high-quality recommendations across all three tasks. We provide a comprehensive overview of the CGR architecture, setting a precedent as the first cybersecurity company to openly discuss these capabilities in such depth. Additionally, we release GUIDE, the largest public collection of real-world security incidents, spanning 13M evidences across 1M annotated incidents.
@@ 1. Introduction
In the rapidly evolving cybersecurity landscape, the sharp rise in threat actors has overwhelmed enterprise security operation centers (SOCs) with an unprecedented volume of incidents to triage (Consulting, 2020). This surge requires solutions that can either partially or fully automate the remediation process. Fully automated systems demand an exceptionally high confidence threshold to ensure correct actions are taken 99% of the time to avoid inadvertently disabling critical enterprise assets. Consequently, attaining such a high level of confidence often renders full automation impractical.
@@
Complexity of security incidents. The extensive variety of security products, each with thousands of custom and built-in detection rules, creates a complex incident landscape further compounded by a scarcity of labeled data.
@@ 1.1. Contributions
We introduce Copilot Guided Response (Fig. 1), an ML framework designed to tackle guided response at scale, marking the first in-depth academic-industry discourse on this vital endeavor. Our framework makes significant contributions to the following areas:
Copilot Guided Response (CGR). The Copilot Guided Response architecture transforms cybersecurity guided response by detailing the first geo-distributed industry-scale framework capable of processing millions of incidents each day with batch latency of just a few minutes. Our ML system scalably deliver three core SOC capabilities—(1) investigation, (2) triaging, and (3) remediation—seamlessly adapting to a range of scenarios, from single alerts to complex incidents involving hundreds of alerts, where each alert is categorized into one of hundreds of thousands of distinct classes, with new classes continuously added.
Largest Cybersecurity Incident Dataset. We introduce GUIDE, the largest publicly available collection of real-world cybersecurity incidents under the permissive CDLA-2.0 license. This extensive dataset includes over 13 million pieces of evidence across 1.6 million alerts and 1M annotated incidents, making it an unparalleled resource for the development and evaluation of GR systems and beyond. By enabling researchers to study real-world data, GUIDE advances the state of cybersecurity and supports the development of next-generation ML systems.
Extensive Evaluation of CGR. We conduct a comprehensive assessment of CGR’s performance across three key pillars: internal evaluations, collaborations with security experts, and feedback from customers. Our internal evaluation on millions of previously unseen incidents show triage models achieving an average cross-region precision of 87% and recall of 41%, with action models reaching 99% precision and 62% recall. Collaboration with Microsoft security research experts confirms the efficacy of our similar incident recommendations, with 94% of incidents deemed relevant, and 98% of security incidents containing one or more recommendations. Furthermore, customer feedback underscores CGR’s overall effectiveness, with 89% of interactions rated positively based confirmation or dismissal rates.
Impact to Microsoft Customers and Beyond. CGR is integrated into the Microsoft Defender XDR product, a leader in the market (Mellen et al., 2024), and is deployed to hundreds of thousands of organizations worldwide. The introduction of CGR to Microsoft Defender XDR has significantly enhanced the operational capabilities of SOCs by streamlining the decision-making process and providing actionable insights across investigation, triaging, and remediation tasks. As a result, Microsoft Defender XDR customers benefit from a more resilient security posture, fortified by adaptive ML-driven guided responses that are tailored to the nuances of their specific security environments.
@@
GR | Guided response (GR) is an ML system that learns from past investigations to generate suggestions |
Copilot | Microsoft Copilot for Security is an AI-powered solution that augments the investigation and response capabilities of security professionals |
To enhance readability, Table 1 details the terminology used in this paper. The reader may want to refer to this table for meanings of technical terms.
@@ 2. Background
We provide an overview of Microsoft Copilot for Security and review literature relevant to guided response.
@@ 2.1. Microsoft Copilot for Security
Microsoft Copilot for Security is an AI-powered solution tailored to enhance the capabilities of security professionals by providing real-time actionable insights and recommendations across Microsoft Defender XDR, Microsoft Sentinel, and Microsoft Intune. At its launch, Copilot for Security introduced five distinct skills: incident summarization, script analyzer, incident report, Kusto query assistant, and guided response (Caparas et al., 2024). The first four skills combine large language models (LLMs), security-specific plugins to contextualize and analyze security data, and post-processing the AI-generated responses to provide key insights (Microsoft, 2024). Guided response is unique by its inclusion of three distinct sub-skills that are not based on natural language—grade recommendation, action recommendation, and similar incident recommendation—which are tailored to the specific preferences of SOCs. Consequently, these sub-skills closely align with traditional machine learning techniques to enable precise and context-specific recommendations.
@@ 2.2. Guided Response
With the market introduction of Microsoft Copilot for Security Guided Response, the concept of guided response was formally defined as “machine learning capabilities to contextualize an incident and learn from previous investigations to generate appropriate response actions” (Gali et al., 2024). Our analysis of academic and industry literature contextualizes relevant contributions within the domain of guided response into three distinct categories: (1) investigation that suggests next steps for further analysis; (2) triaging to determine whether an incident is a true positive, false positive, or benign positive (informational); and (3) remediation which proposes specific response actions to contain and resolve incidents. While categorizing guided response into these three areas, our analysis highlights the multifaceted role of machine learning in modern cybersecurity.
@@
Investigation. Assisting security analysts in their investigation of incidents is a pivotal aspect of cybersecurity. Machine learning assisted investigation typically encompasses: (1) similar incident identification (Zhong et al., 2018; Jiang et al., 2024; Lin, 2018), (2) investigation assistance (Franco et al., 2020; Nguyen and Choo, 2021; Perera et al., 2019), and (3) playbook recommendation (Kremer et al., 2023; Kraeva and Yakhyaeva, 2021; Applebaum et al., 2018). Our research focuses on similar incident recommendation, an understudied topic in the context of SOCs. While a few industry solutions offer similar incident identification capabilities through various methods, such as comparing and counting identical artifacts (Networks, 2024b, a), details regarding their methodologies and performance metrics are scarce. We address this gap by detailing the first industry-scale architecture for similar incident recommendation in SOCs.
@@
Triaging. Incident triaging is a vital and time intensive initial task typically performed by junior security analysts who sort through large volumes of incidents to identify ones that require further investigation. This process includes prioritizing incidents for in-depth investigation (Oprea et al., 2018; Aminanto et al., 2020; Liu et al., 2022; Alsubhi et al., 2012, 2008), and filtering them based on the likelihood of being true positive versus benign or false positive (Ban et al., 2021; Sopan et al., 2018; Hassan et al., 2019, 2020). Several leading security industry solutions employ ML to automate the triaging process by utilizing threat intelligence feeds, correlated alerts, or historical data (IBM, 2020). However, there remains a significant lack of public information regarding their architectures or performance. While some security companies provide insights into their triaging systems (Sopan et al., 2018; Feng et al., 2017), the evaluations take place under constrained scenarios and lack critical deployment details. We address these gaps by detailing an ML triaging architecture that can scale, adapt, and deploy to hundreds of thousands of SOCs around the world.
@@
Remediation. The majority of research related to SOC remediation occurs within the domain of intrusion response systems (IRS). IRS are designed to notify SOC analysts, or in certain cases, dynamically respond to detected intrusions. These systems leverage various decision-making models, including rule-based approaches (Foo et al., 2005), multi-objective optimization (Li et al., 2018; Shameli-Sendi et al., 2016; Toth and Kruegel, 2002), game theory (Huang et al., 2015; Qin et al., 2018; Zonouz et al., 2013), reinforcement learning (Alturkistani and El-Affendi, 2022; Stefanova and Ramachandran, 2018; Khoury and Nassar, 2020), and many more (Bashendy et al., 2023; Inayat et al., 2016; Shameli-Sendi et al., 2012) to select optimal responses based on the current state of the system, the nature of the attack, and the potential impact of the countermeasures. Despite the advancements in IRS, there is limited discussion on how to scale these systems to handle industry demands, which includes managing millions of incidents per day, integrating with complex incident scenarios involving hundreds of alerts across hundreds of thousands of classes, and customizing responses to specific SOC preferences. While a few industry solutions employ human-in-the-loop IRS to aid in the remediation process, these are primarily focused on endpoint threats and provide static remediation suggestions (Networks, 2024c). Our research aims to bridge this divide, enhancing both the scalability and transparency of industry-scale guided remediation systems.
@@ 3. Architecture Overview
We overview the Copilot Guided Response (CGR) architecture, organized around three key pipelines: train, inference, and embedding, as illustrated in Figure 1. To manage heavy data preprocessing efficiently, we utilize PySpark’s distributed computational engine whenever possible, while reserving Python for last mile recommendation tasks that do not have PySpark support. Below, we detail how these pipelines synergize to guide security analysts through the processes of investigating, triaging, and remediating security incidents across the vast enterprise landscape.
@@
To adhere to privacy regulations, CGR is uniformly replicated across each geographic region, utilizing Synapse to ensure consistency and compliance. Consequently, our discussion in the following sections focuses on the development of CGR from the perspective of a single geographic region.
@@ 4.1. Preprocessing
T1—Feature engineering. We collect alert telemetry from multiple Azure Data Lake Storage (ADLS) tables and join them into a PySpark alert dataframe. Each row in the alert dataframe contains columns for unique alert and incident identifiers, complemented by customer-provided grade and remediation action, when available. Additionally, each row contains 5 categorical feature columns—OrganizationId, DetectorId, ProductId, Category, and Severity—along with 67 engineered numerical feature columns, developed in close collaboration with Microsoft security research experts. We retain rows even if they lack a customer grade or action, as these alerts can merge with other alerts to form incidents that do contain such data.
T2—Feature space compression. Before converting the categorical columns to one-hot-encoded representations, we must address the challenge of high cardinality in the DetectorId and OrgId columns. In various geographic regions, DetectorIds can exceed 100,000 and OrgIds can reach up to 50,000, creating an extremely large and sparse feature space that often leads to failures during dimensionality reduction in the PySpark Synapse cluster. To mitigate this, we aggregate the feature space by substituting infrequent values—those associated with fewer than 10 alerts—with a generic value. This method, while resulting in some information loss, ensures the system remains within the computational boundaries of the PySpark cluster.
@@
T6—Dimensionality reduction. We independently apply principal component analysis (PCA) to both the incident and alert dataframes, each containing tens of thousands of columns. Leveraging PySpark’s native machine learning library and its distributed computing capabilities allows us to efficiently reduce the feature space, preventing out-of-memory errors when centralizing the dataframes to the primary PySpark node for scikit-learn model training. Our goal is to reduce the feature space into principal components that captures 95% of the original variance in each dataframe; our findings indicate that setting to 40 achieves this goal. The resulting PCA model weights are stored in an Azure Blob Storage container for use in the inference and embedding pipelines.
T7—Store embeddings. The final step is to save the incident embeddings to an ADLS table to enhance the similar incident recommendation algorithm within the inference pipeline. A strategic product decision dictates that only the top five most similar incidents are displayed for any given incident. Consequently, we only store up to five instances of each incident, categorized by triage grade and the unique set of DetectorIds that comprise the IncidentHash.
@@ 4.2. Model Training
T9—Training process. We select a random forest model due to its efficiency on our CPU-based PySpark infrastructure and its reliable performance with tabular data. We conduct a grid search over four key model parameters: , , , and , and select the model with the highest macro-F1 score on the validation set.
T10—Validation and model storage. Once the best model has been identified for both the triage and remediation tasks, we compare them with previous models to ensure quality between training cycles remains within a few percentage points. We do not require each model to have a higher macro-F1 score since new detectors and security products are onboarded over time, which can cause fluctuations in performance. After validation, the models are saved to an Azure Blob Storage container for use in the inference pipeline.
@@ 5. Inference Pipeline
Leveraging the infrastructure outlined in Section 4, the inference pipeline begins with the preprocessing of batched alert data to facilitate guided response recommendations (Sec 5.1). The trained models and historical embeddings are then employed across three sequential phases: triage to predict incident grades (Sec 5.2), investigation to identify similar incidents (Sec 5.3), and remediation to determine response actions (Sec 5.4). As incidents evolve, recommendations are dynamically updated to incorporate the latest information. Upon completion, the inference pipeline stores these guided response recommendations in a table, ensuring rapid access for customers using the Microsoft Defender XDR product. A step-by-step overview of the entire inference process is outlined in Algorithm 2.
@@ 5.1. Preprocessing
The initial phase of the inference pipeline prepares real-time batched alert data for guided response recommendations. The process begins by retrieving the last 15 minutes of alert telemetry and loading them into a PySpark dataframe. These alerts are then processed using the feature space compression and one-hot encoding techniques outlined in Section 4.1. Following this, we bifurcate the alert data into two distinct PySpark dataframes—one dedicated to generating remediation predictions and the other for aggregating alerts into incidents for similar incident and triage recommendations, utilizing the aggregation process described in Section 4.1. Next, we apply the latest PCA models from the training pipeline to reduce the dimensionality of both dataframes to form alert and incident embeddings. The incident embeddings are then stored in an ADLS table to enhance the similar incident recommendations, with a limit of five instances per incident, categorized by triage grade and DetectorId. Finally, we convert both PySpark dataframes into Pandas dataframes to facilitate subsequent triage, investigation, and remediation processes.

### v4

@@ Abstract.
Security operation centers contend with a constant stream of security incidents, ranging from straightforward to highly complex. To address this, we developed Microsoft Copilot for Security Guided Response (CGR), an industry-scale ML architecture that guides security analysts across three key tasks—(1) investigation, providing essential historical context by identifying similar incidents; (2) triaging to ascertain the nature of the incident—whether it is a true positive, false positive, or benign positive; and (3) remediation, recommending tailored containment actions. CGR is integrated into the Microsoft Defender XDR product and deployed worldwide, generating millions of recommendations across thousands of customers. Our extensive evaluation, incorporating internal evaluation, collaboration with security experts, and customer feedback, demonstrates that CGR delivers high-quality recommendations across all three tasks. We provide a comprehensive overview of the CGR architecture, setting a precedent as the first cybersecurity company to openly discuss these capabilities in such depth. Additionally, we release GUIDE, the largest public collection of real-world security incidents, spanning 13M evidences across 1M incidents annotated with ground-truth triage labels by customer security analysts. This dataset represents the first large-scale cybersecurity resource of its kind, supporting the development and evaluation of guided response systems and beyond.
@@ 1. Introduction
In the rapidly evolving cybersecurity landscape, the sharp rise in threat actors has overwhelmed enterprise security operation centers (SOCs) with an unprecedented volume of incidents to triage (Consulting, 2020). This surge requires solutions that can either partially or fully automate the remediation process. Fully automated systems demand an exceptionally high confidence threshold (e.g., 99%) to ensure correct actions are taken to avoid inadvertently disabling critical enterprise assets. Consequently, attaining such a high level of confidence often renders full automation impractical.
@@
Complexity of security incidents. The extensive variety of security products, each with thousands of custom and built-in detectors, creates a complex incident landscape further compounded by a scarcity of labeled data.
@@ 1.1. Contributions
We introduce Copilot Guided Response (Fig. 1), an ML framework designed to tackle guided response at scale. Our framework makes significant contributions in the following areas:
Copilot Guided Response (CGR). The Copilot Guided Response architecture transforms cybersecurity guided response by detailing the first geo-distributed industry-scale framework capable of processing millions of incidents each day with batch latency of just a few minutes. Our ML system scalably delivers three core SOC capabilities—(1) investigation, (2) triaging, and (3) remediation—seamlessly adapting to a range of scenarios, from single alerts to complex incidents involving hundreds of alerts, where each alert is categorized into one of hundreds of thousands of distinct classes, with new classes continuously added.
Largest Cybersecurity Incident Dataset. We introduce GUIDE, the largest publicly available collection of real-world cybersecurity incidents under the permissive CDLA-2.0 license. This extensive dataset includes over 13 million pieces of evidence across 1.6 million alerts and 1 million incidents annotated with ground-truth triage labels by customer security analysts, making it an unparalleled resource for the development and evaluation of GR systems and beyond. By enabling researchers to study real-world data, GUIDE advances the state of cybersecurity and supports the development of next-generation ML systems.
Extensive Evaluation of CGR. We provide a comprehensive evaluation of CGR’s performance, focusing on its industrial applicability rather than comparisons with alternative designs or competing security products, which are often undisclosed. The release of the GUIDE dataset enables researchers to explore new architectures for maximum performance. Our evaluation spans internal testing, expert collaborations, and customer feedback. Internal assessments on hundreds of thousands of unseen incidents show triage models achieve 87% precision and 41% recall, while action models reach 99% precision and 62% recall. Collaboration with Microsoft security experts confirms the efficacy of our similar incident recommendations, with 94% of incidents deemed relevant, and 98% of incidents containing one or more recommendations. Customer feedback further underscores its effectiveness, with 89% of interactions rated positively.
Impact to Microsoft Customers and Beyond. CGR is integrated into the Microsoft Defender XDR product, a leader in the market (Mellen et al., 2024), and is deployed to hundreds of thousands of organizations worldwide. The introduction of CGR to Microsoft Defender XDR has significantly enhanced the operational capabilities of SOCs by streamlining the decision-making process and providing actionable insights across investigation, triaging, and remediation tasks. As a result, Microsoft Defender XDR customers benefit from a more resilient security response, fortified by adaptive ML-driven guided responses that are tailored to the nuances of their specific security environments.
@@ 2. Background
We provide an overview of Microsoft Copilot for Security and review literature relevant to guided response. To enhance readability, Table 1 details the terminology used in this paper.
@@ 2.1. Microsoft Copilot for Security
Microsoft Copilot for Security is an AI-driven solution that enhances security professionals’ workflows by offering real-time insights and recommendations across Microsoft Defender XDR, Microsoft Sentinel, and Microsoft Intune. At launch, it introduced five skills: incident summarization, script analyzer, incident report, Kusto query assistant, and guided response (Caparas et al., 2024). While the first four leverage LLMs with security-specific plugins and post-processing (Microsoft, 2024), guided response includes three machine learning sub-skills—grade recommendation, action recommendation, and similar incident recommendation—tailored to SOC preferences for precise, context-specific insights.
@@ 2.2. Guided Response
With the market introduction of Microsoft Copilot for Security Guided Response, the concept of guided response was formally defined as “machine learning capabilities to contextualize an incident and learn from previous investigations to generate appropriate response actions” (Gali et al., 2024). Our analysis of academic and industry literature contextualizes relevant contributions within the domain of guided response into three distinct categories: (1) investigation that suggests next steps for further analysis; (2) triaging to determine whether an incident is a true positive, false positive, or benign positive (e.g., informational); and (3) remediation which proposes specific response actions to contain and resolve incidents.
@@
Investigation. Assisting security analysts in their investigation of incidents is a pivotal aspect of cybersecurity. Machine learning assisted investigation typically encompasses: (1) similar incident identification (Zhong et al., 2018; Jiang et al., 2024; Lin, 2018), (2) investigation assistance (Franco et al., 2020; Nguyen and Choo, 2021; Perera et al., 2019), and (3) playbook recommendation (Kremer et al., 2023; Kraeva and Yakhyaeva, 2021; Applebaum et al., 2018). Our research focuses on similar incident recommendation, an understudied topic in the context of SOCs. While a few industry solutions offer similar incident identification capabilities through methods such as comparing and counting identical artifacts (Networks, 2024b, a), details regarding their methodologies and performance metrics are scarce. We address this gap by detailing the first industry-scale architecture for similar incident recommendation in SOCs.
@@
Triaging. Incident triaging is a vital and time intensive task typically performed by junior analysts to identify incidents requiring further investigation. This involves prioritizing incidents for deeper review (Oprea et al., 2018; Aminanto et al., 2020; Liu et al., 2022; Alsubhi et al., 2012, 2008) and filtering them based on the likelihood of being true versus false positives (Ban et al., 2021; Sopan et al., 2018; Hassan et al., 2019, 2020). While some industry solutions use ML to automate triaging (IBM, 2020), there is limited public information on their architecture or performance. Although some companies provide insights (Sopan et al., 2018; Feng et al., 2017; Oliver et al., 2024), these are often limited to controlled scenarios and lack critical deployment details. Our work addresses these gaps by detailing a scalable, adaptable ML triaging architecture that is deployed globally to thousands of SOCs.
@@
Remediation. The majority of remediation research centers on intrusion response systems (IRS), designed to notify SOC analysts or dynamically respond to detected intrusions. These systems employ various decision-making models, including rule-based approaches (Foo et al., 2005), multi-objective optimization (Li et al., 2018; Shameli-Sendi et al., 2016; Toth and Kruegel, 2002), game theory (Huang et al., 2015; Qin et al., 2018; Zonouz et al., 2013), human-in-the-loop (Networks, 2024c), reinforcement learning (Alturkistani and El-Affendi, 2022; Stefanova and Ramachandran, 2018; Khoury and Nassar, 2020), and many others (Bashendy et al., 2023; Inayat et al., 2016; Shameli-Sendi et al., 2012) to determine optimal responses based on the system state, nature of the attack, and countermeasure impact. Despite advancements, there is limited discussion on how to scale these systems to handle industry demands, such as managing millions of incidents, handling complex scenarios with numerous alerts, and customizing responses for SOC preferences. Our research bridges this gap, enhancing the scalability and transparency of industry-scale guided remediation systems.
@@ 3. Architecture Overview
We detail the Copilot Guided Response (CGR) architecture, organized around three key pipelines: train, inference, and embedding, as illustrated in Figure 1. We utilize PySpark’s distributed computational engine whenever possible, while reserving Python for last mile recommendation tasks that do not have PySpark support. Below, we detail how these pipelines synergize to guide security analysts through the processes of investigating, triaging, and remediating security incidents across the enterprise landscape.
@@
To adhere to privacy regulations, CGR is replicated across geographic regions, utilizing Synapse to ensure consistency and compliance. Consequently, the following sections focuses on development from the perspective of a single geographic region.
@@ 4.1. Preprocessing
T1—Feature engineering. We collect alert telemetry from multiple Azure Data Lake Storage (ADLS) tables and join them into a PySpark alert dataframe. Each row in the alert dataframe contains columns for unique alert and incident identifiers, complemented by customer-provided grade and remediation action, when available. Additionally, each row contains 5 categorical feature columns—OrganizationId, DetectorId, ProductId, Category, and Severity—along with 67 engineered numerical feature columns, developed in close collaboration with Microsoft security research experts. We retain rows that lack a customer grade or action, as these alerts can merge with other alerts to form incidents that do contain labels.
T2—Feature space compression. Before converting the categorical columns to one-hot-encoded representations, we must address the challenge of high cardinality in the DetectorId and OrgId columns. In various geographic regions, DetectorIds can exceed 100k and OrgIds can reach up to 50k, creating an extremely large and sparse feature space that often leads to failures during dimensionality reduction in the PySpark cluster. To mitigate this, we aggregate the feature space by substituting infrequent values—those associated with fewer than 10 alerts—with a generic value. This method, while resulting in some information loss, ensures the system remains within the computational boundaries of the cluster.
@@
T6—Dimensionality reduction. We independently apply principal component analysis (PCA) to both the incident and alert dataframes, each containing tens of thousands of columns. PCA is chosen for its performane and availability within PySpark’s native machine learning library, which supports distributed computing and allows for efficient feature space reduction. This mitigates the risk of out-of-memory errors that occur when centralizing large dataframes on the primary PySpark node for subsequent scikit-learn model training. Our objective is to condense the feature space to principal components that captures 95% of the original variance in each dataframe. Empirically, we find that setting meets this requirement. The resulting PCA weights are saved in an Azure Blob Storage container, enabling reuse within the inference and embedding pipelines.
T7—Store embeddings. The final step is to save the incident embeddings to an ADLS table to enhance similar incident recommendations within the inference pipeline. A strategic product decision dictates that only the top five most similar incidents are displayed for any given incident. Consequently, we only store up to five instances of each incident, categorized by triage grade and the unique set of DetectorIds that comprise the IncidentHash.
@@ 4.2. Model Training
T9—Training process. We select a random forest model due to its efficiency on our CPU-based PySpark infrastructure and its reliable performance on tabular data. We conduct a grid search over four key model parameters: , , , and , and select the model with the highest macro-F1 score on the validation set.
T10—Validation and model storage. Once the best model has been identified for both triage and remediation tasks, we compare them with previous models to ensure quality between training cycles remains within a few percentage points. We do not require each model to have a higher macro-F1 score since new detectors and security products are onboarded over time, which can cause fluctuations in performance. After validation, the models are saved to an Azure Blob Storage container for use in the inference pipeline.
@@ 5. Inference Pipeline
Building on the infrastructure in Section 4, the inference pipeline processes batched data to provide guided response recommendations (Sec 5.1). For each new or updated incident, the trained models and historical embeddings operate across three phases: triage to predict grades (Sec 5.2), investigation to find similar incidents (Sec 5.3), and remediation to recommend response actions (Sec 5.4). Recommendations dynamically update as incidents evolve and are stored in a table, ensuring rapid access for customers. Algorithm 1 provides a step-by-step overview.
@@ 5.1. Preprocessing
The initial phase of the inference pipeline prepares real-time batched alert data, retrieving the last 15 minutes of telemetry and loading it into a PySpark dataframe. These alerts are then processed using the feature space compression and one-hot encoding techniques outlined in Section 4.1. Next, we bifurcate the alert data into two distinct PySpark dataframes—one dedicated to generating remediation predictions and the other for aggregating alerts into incidents for similar incident and triage recommendations, utilizing the aggregation process described in Section 4.1. Afterwards, we apply the latest PCA models from the training pipeline to reduce the dimensionality of both dataframes to form alert and incident embeddings. The incident embeddings are stored in an ADLS table to enhance the similar incident recommendations, with a limit of five instances per incident, categorized by triage grade and DetectorId. Finally, we convert both PySpark dataframes into Pandas dataframes to facilitate subsequent triage, investigation, and remediation processes.